# DynamicControl : Adaptive Condition Selection for Improved Text-to-Image Generation

### v1

@@ Abstract
To enhance the controllability of text-to-image diffusion models, current ControlNet-like models have explored various control signals to dictate image attributes. However, existing methods either handle conditions inefficiently or use a fixed number of conditions, which does not fully address the complexity of multiple conditions and their potential conflicts. This underscores the need for innovative approaches to manage multiple conditions effectively for more reliable and detailed image synthesis. To address this issue, we propose a novel framework, DynamicControl , which supports dynamic combinations of diverse control signals, allowing adaptive selection of different numbers and types of conditions. Our approach begins with a double-cycle controller that generates an initial real score sorting for all input conditions by leveraging pre-trained conditional generation models and discriminative models. This controller evaluates the similarity between extracted conditions and input conditions, as well as the pixel-level similarity with the source image. Then, we integrate a Multimodal Large Language Model (MLLM) to build an efficient condition evaluator. This evaluator optimizes the ordering of conditions based on the double-cycle controller’s score ranking. Our method jointly optimizes MLLMs and diffusion models, utilizing MLLMs’ reasoning capabilities to facilitate multi-condition text-to-image (T2I) tasks. The final sorted conditions are fed into a parallel multi-control adapter, which learns feature maps from dynamic visual conditions and integrates them to modulate ControlNet, thereby enhancing control over generated images. Through both quantitative and qualitative comparisons, DynamicControl demonstrates its superiority over existing methods in terms of controllability, generation quality and composability under various conditional controls.
@@ 3.3 Condition Evaluator
$R{}$ $=LLaVA(c_{1},c_{2},...,c_{N},\tau^{\prime};\omega)$ (7)
$h_{i}$ $=H(c_{1},c_{2},...,c_{N},\tau^{\prime};\omega|r_{i})$
$f_{c}$ $=Q(\mathcal{H})$
@@
$\mathcal{L}_{\mathrm{LLM}}(\tau)=-\sum_{i=1}^{N}\log p_{\omega+\Delta\omega(\theta)}(\textless con^{i}\textgreater\mid c_{1},c_{2},...,c_{N},\tau^{\prime})$ (8)
@@
$\mathcal{L}_{\text{eval}}=-\sum_{i=1}^{N}c_{i}\log p_{i}$ (9)
@@ 3.5 Training Strategy
The whole training of our network consists of two processes. The first process involves training the condition evaluator. The overall loss function, which combines LLM and diffusion losses, is represented as follows::
$\mathcal{L}_{\text{condi}}=\lambda_{1}\mathcal{L}_{\text{condition}}+\lambda_{2}\mathcal{L}_{\text{image}}+\lambda_{3}\mathcal{L}_{\text{LLM}}+\lambda_{4}\mathcal{L}_{\text{eval}}$ (10)
where $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$ and $\lambda_{4}$ are positive constants to balance the different losses, which are set to 1, 1, 2 and 1.5 respectively in our practice. This evaluator is then frozen, remaining unchanged during any subsequent optimization processes.
@@ 4 Experiments
We validate the effectiveness of DynamicControl on five conditions with more common control conditions: canny, hed, segmentation mask, openpose and depth. Our evaluation primarily focuses on several leading methods in the realm of controllable text-to-image diffusion models, including Gligen [29], T2I-Adapter [35], ControlNet v1.1 [63], GLIGEN [29], Uni-ControlNet [65], UniControl [39], Cocktail [16] and ControlNet++ [27]. These methods are pioneering in their field and provide public access to their codes and model weights, which accommodate various image conditions. Although the models of other approaches such as AnyControl [54] are public, their code cannot be successfully run after many attempts. Implementation details including network structure, datasets, evaluation metrics, hyper-parameters of training and inference can be found in the supplementary material.
@@ 4.2 Qualitative Comparison
In Fig. 5 and Fig. A3, we visually compare different tasks — Canny, HED, Depth, Segmentation, Openpose, Normal, Hedsketch conditions. Our method consistently outperforms other models in terms of both visual quality and alignment with the specified conditions or prompts. This demonstrates the effectiveness of our approach in handling a diverse range of image generation tasks, while maintaining high fidelity to the input conditions. For instance, for the Hed and Segmentation task, the results generated by our model exhibit a higher degree of detail preservation and visual consistency. The outputs of our method maintain a faithful reproduction of the edge or mask information compared to other methods. In the Depth task, our model exhibits its superiority by producing plausible images characterized by seamless transitions and textures that mimic natural appearances. Furthermore, our model exhibits a more nuanced comprehension of geometric guidance pertaining to openpose and surface normals than the others. The outputs conditioned on openpose are noticeably more precise. The comparative visual analysis highlights the robustness and adaptability of our method, demonstrating its efficacy across a broad spectrum of tasks.
@@ 5 Conclusion
Training Details. During the first stage of training, we adopt the pre-trained LLaVAv1.1-7B [32] and QFormer [26] and employ DeepSpeed [1] Zero-2 to perform LoRA [15] fine-tuning. The Stable Diffusion-1.5 [44] is diffusion model pre-trained weights. The learning rate and weight decay parameters are set to 2e-4 and 0, respectively. In the second stage, the values of learning rate, weight decay, and warm-up ratio are set to 1e-5, 0, and 0.001, respectively. We train the model for totally 50K iterations. Furthermore, we take the AdamW [33] as the optimizer based on PyTorch Lightning [37] for both stages. Our full-version DynamicControl is trained on 48 Nvidia-V100 GPUs with the batch size of 4. In all the experiments, we adopt DDIM [51] sampler with 50 timesteps for all the compared methods.

### v2

@@ Abstract
To enhance the controllability of text-to-image diffusion models, current ControlNet-like models have explored various control signals to dictate image attributes. However, existing methods either handle conditions inefficiently or use a fixed number of conditions, which does not fully address the complexity of multiple conditions and their potential conflicts. This underscores the need for innovative approaches to manage multiple conditions effectively for more reliable and detailed image synthesis. To address this issue, we propose a novel framework, DynamicControl , which supports dynamic combinations of diverse control signals, allowing adaptive selection of different numbers and types of conditions. Our approach begins with a double-cycle controller that generates an initial real score sorting for all input conditions by leveraging pre-trained conditional generation models and discriminative models. This controller evaluates the similarity between extracted conditions and input conditions, as well as the pixel-level similarity with the source image. Then, we integrate a Multimodal Large Language Model (MLLM) to build an efficient condition evaluator. This evaluator optimizes the ordering of conditions based on the double-cycle controller’s score ranking. Our method jointly optimizes MLLMs and diffusion models, utilizing MLLMs’ reasoning capabilities to facilitate multi-condition text-to-image (T2I) tasks. The final sorted conditions are fed into a parallel multi-control adapter, which learns feature maps from dynamic visual conditions and integrates them to modulate ControlNet, thereby enhancing control over generated images. Through both quantitative and qualitative comparisons, DynamicControl demonstrates its superiority over existing methods in terms of controllability, generation quality and composability under various conditional controls. The code and models will be available for further research.
@@ 3.3 Condition Evaluator
$R{}$ $=LLaVA(c_{1},c_{2},...,c_{N},\tau^{\prime};\omega),$ (7)
$h_{i}$ $=H(c_{1},c_{2},...,c_{N},\tau^{\prime};\omega|r_{i}),$
$f_{c}$ $=Q(\mathcal{H}),$
@@
$\mathcal{L}_{\mathrm{LLM}}(\tau)=-\sum_{i=1}^{N}\log p_{\omega+\Delta\omega(\theta)}(\textless con^{i}\textgreater\mid c_{1},c_{2},...,c_{N},\tau^{\prime}).$ (8)
@@
$\mathcal{L}_{\text{eval}}=-\sum_{i=1}^{N}c_{i}\log p_{i}.$ (9)
@@ 3.5 Training Strategy
The whole training of our network consists of two processes. The first process involves training the condition evaluator. The overall loss function, which combines LLM and diffusion losses, is represented as follows:
$\mathcal{L}_{\text{condi}}=\lambda_{1}\mathcal{L}_{\text{condition}}+\lambda_{2}\mathcal{L}_{\text{image}}+\lambda_{3}\mathcal{L}_{\text{LLM}}+\lambda_{4}\mathcal{L}_{\text{eval}},$ (10)
where $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$ and $\lambda_{4}$ are positive constants to balance the different losses. This evaluator is then frozen, remaining unchanged during any subsequent optimization processes.
@@ 4 Experiments
We validate the effectiveness of DynamicControl on five conditions with more common control conditions: canny, hed, segmentation mask, openpose and depth. Our evaluation primarily focuses on several leading methods in the realm of controllable text-to-image diffusion models, including Gligen [29], T2I-Adapter [35], ControlNet v1.1 [63], GLIGEN [29], Uni-ControlNet [65], UniControl [39], Cocktail [16] and ControlNet++ [27]. These methods are pioneering in their field and provide public access to their codes and model weights, which accommodate various image conditions. Although the models of other approaches such as AnyControl [54] are public, their code cannot be successfully run after many attempts. Implementation details including network structure, datasets, evaluation metrics, computational complexities, hyper-parameters of training and inference can be found in the supplementary material.
@@ 4.2 Qualitative Comparison
In Fig. 5 and Fig. 6, we visually compare different tasks — Canny, HED, Depth, Segmentation, Openpose, Normal, Hedsketch conditions. Our method consistently outperforms other models in terms of both visual quality and alignment with the specified conditions or prompts. This demonstrates the effectiveness of our approach in handling a diverse range of image generation tasks, while maintaining high fidelity to the input conditions. For instance, for the Hed and Segmentation task, the results generated by our model exhibit a higher degree of detail preservation and visual consistency. The outputs of our method maintain a faithful reproduction of the edge or mask information compared to other methods. In the Depth task, our model exhibits its superiority by producing plausible images characterized by seamless transitions and textures that mimic natural appearances. Furthermore, our model exhibits a more nuanced comprehension of geometric guidance pertaining to openpose and surface normals than the others. The outputs conditioned on openpose are noticeably more precise. The comparative visual analysis highlights the robustness and adaptability of our method, demonstrating its efficacy across a broad spectrum of tasks.
@@ 5 Conclusion
Training Details. During the first stage of training, we adopt the pre-trained LLaVAv1.1-7B [32] and QFormer [26] and employ DeepSpeed [1] Zero-2 to perform LoRA [15] fine-tuning. The Stable Diffusion-1.5 [44] is diffusion model pre-trained weights. The learning rate and weight decay parameters are set to 2e-4 and 0, respectively. In the second stage, the values of learning rate, weight decay, and warm-up ratio are set to 1e-5, 0, and 0.001, respectively. We train the model for totally 50K iterations. Furthermore, we take the AdamW [33] as the optimizer based on PyTorch Lightning [37] for both stages. Our full-version DynamicControl is trained on 8 H20 GPUs with the batch size of 4 and the whole training process can be completed within 3 days. It takes an average of 34s to generate one image on one H20 GPU. The values of $\lambda$ in Eq.10 are experimentally tested through a series of experiments, which are set to 1, 1, 2 and 1.5 respectively in our practice. These values vary from 0.1 to 3 at every 0.1 interval. In all the experiments, we adopt DDIM [51] sampler with 50 timesteps for all the compared methods.
@@
A2.1 Comparison on computational complexities
As shown in Tab. A1, we report the GPU memory and inference time for one image comparison results of different methods. It is worth noting that compared with other methods, our method is still in the normal range in terms of resource consumption and inference time although we use MLLM.
Ours ControlNet++ Uni-ControlNet Uni-Control T2I-Adapter Cocktail
40G / 34s 48G / 30s 32G / 33s 40G / 39s 32G / 43s 80G / 55s
Table A1: GPU memory(G) / inference time(s) comparison of different methods.
A2.2 Generation to mismatched conditions
As shown in Fig. A2, we illustrate the results of multiple mismatched conditions, which demonstrates that based on the archetype and data, our method has learned the combination of different conditions. But we still want to clarify that this non-aligned control generation is completely different from our setting. What we want to address is the complexity of multiple conditions and their potential conflicts for the same subject as we describe in the abstract and introduction. And we will dive into the exploration of this non-aligned control generation task in the future.
Figure A2: Results of multiple conditions correspond to different subjects.
A2.3 User Study